The Shocking Corporation is subject to a company tax rate of 30%.

The company asks you to calculate the after tax cost of each of the following sources of funds:

(a) Ordinary $1 shares expect the next dividend payment to be $0.102 per share which is expected to grow at a rate of 2% in perpetuity. The current market price is$ 1.20.

(b) What is the cost of $100,000 in retained earnings?

(c) 10% preference shares of $2 with a current market price of $2.50.

(d) 9% debentures with an issued value of $100,redeemable in four years at $100 with a current market price of $90.90.
Answer #1

A:

Price of equity share= Dividend expected/ (Cost of equity shares- Growth rate)

1.2 = 0.102/ (Cost of equity shares- 0.02)

Cost of equity shares = 0.102/1.2 + 0.02

=10.5%

B: Cost of retained earnings = Cost of equity shares (Since there is no flotation cost)

= 10.5%

C: Cost of preference shares = Dividend/ Price

= 10%*2/ 2.5

=8%

D: Cost of debentures = Rd*(1-Tax)

Rd is YTM computed using financial calculator

Input: PMT = 9%*100 = 9

FV = 100

PV= 90.90

N= 4

Find I/Y = 12%

Cost of debentures = 12%*(1-0.3) = 8.4%